# On the growth of actions of free products

### v1

@@ 3.1. A criterion
Let $G$ be a finitely generated group endowed with a finite symmetric generating set $S$. If $X$ is a $G$-set, we endow $X$ with the its Schreier graph structure. The associated simplicial distance is denoted $d_{G,X}$. We denote by $B_{G,X}(x,n)$ the corresponding balls. The support of an element $g\in G$ in $X$ is the set $\textrm{supp}_{X}(g)=\{x\in X\colon gx\neq x\}$. Given $R>0$, we shall say that two subsets $A,B\subset X$ are $R$-separated if $d_{G,X}(a,b)\geq R$ for every $a\in A$ and $b\in B$. Finally the $R$-coarse connected components of a subset $A\subset X$ are the equivalence classes of the equivalence relation on $A$ generated by the pairs $(a_{1},a_{2})$ such that $d_{G,X}(a_{1},a_{2})\leq R$.
@@
Let $G$ be a finitely generated group endowed with a finite symmetric generating set $S$. Let $\alpha,\beta\colon\mathbb{N}\to\mathbb{N}$ be functions such that $\alpha(n)\preccurlyeq\beta(n)$. We say that $G$ satisfies the sparse support condition at scale $(\alpha,\beta)$, if for every faithful $G$-set $X$ such that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\beta(n)$, there exists constants $C,D>0$ such that for every $R>0$, there exists a non-trivial element $g\in G$ satisfying the following
@@
This condition does not depend on the choice of a finite generating set $S$ for $G$. Note that condition (C4) implies in particular that $G$ satisfies a Schreier growth gap $\alpha(n)$. The sparse support condition implies that this gap can be improved for any non-trivial free product with $G$, as follows.
@@
Set $L=G\ast H$. We fix a finite generating set of $L$ of the form $S\cup T$, where $S$ is a generating set of $G$ and $H$ is a generating set of $H$. Let $t\in T$ be a non-trivial. Let $X$ be a faithful $L$-set. We shall consider on $X$ the Schreier graph distance by $d_{L,X}$, as well as the distance $d_{G,X}$ induced by restricting the action to $G$. Suppose that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\beta(n)$ (else, the desired conclusion is true trivially). Let $C,D$ be as in Definition 3.1, fix $R>0$ (which we may assume is even), and let $g\in G$ be the corresponding element. Consider the commutator $h=[g,t]$. Note that $\textrm{supp}_{X}(h)$ is contained in the 1-neighbourhood of $\textrm{supp}_{X}(g)$ (with respect to the distance $d_{L,X}$).
Since $h$ has infinite order and the action is faithful, we can find $x_{0}\in X$ such that the points $x_{n}=h^{n}(x_{0})$ are pairwise distinct for $n=0,\ldots,R$. Note that $d_{L,X}(x_{n},x_{n+1})\leq 2D+2$, by condition (C1). For each point $x_{n}$ we choose $y_{n}\in\textrm{supp}_{X}(g)$ such that $d_{L,X}(x_{n},y_{n})\leq 1$.. Then all points $y_{0},\ldots,y_{R}$ belong to the ball $B_{L,X}(x_{0},D_{1}R)$, with $D_{1}=2D+3$. Since all $D$-coarse connected components of $\textrm{supp}_{X}(g)$ with respect the distance $d_{G,X}$ have diameter bounded by $C$, their 1-neighbourhood in the distance $d_{L,X}$ has cardinality uniformly bounded by some constant $C_{1}>0$ (not depending on $R$).
@@ 3.2. Confined subgroups
Equivalently, a subgroup $H$ is confined if the closure of the set of conjugates of $H$ in the space $\operatorname{Sub}(G)$ of subgroups of $G$, does not contain the trivial subgroup. The following simple lemma explains the usefulness of the notion of confined subgroups for the study of growth of actions. See [LBMB22a, Lemma 1.8] for a proof.
@@
We will invoke the following, which is a particular case of Proposition 1.6 in [LBMB22a].
@@ 3.3. Houghton groups
If $\Omega$ is a set, $\operatorname{Sym}(\Omega)$ is the group of all self-bijections of $\Omega$. We denote by $\operatorname{Alt}_{f}(\Omega)$ the group of alternating finitely supported permutations of $\Omega$. For $H\leq\operatorname{Sym}(\Omega)$, write $\Omega_{H,f}$ for the union of finite orbits under $H$, and $\Omega_{H,\infty}$ for the union of infinite ones. We say that a partition is $H$-invariant if every element of $H$ sends a block of the partition to a (possibly different) block of the partition.
The following is a special case of [LBMB22b, Proposition 3.7].
@@
Let $\Omega$ be a countable set, and let $G\leq\operatorname{Sym}(\Omega)$ be any subgroup containing $\operatorname{Alt}_{f}(\Omega)$. Then a subgroup $H\leq G$ is confined if and only if $\Omega_{H,f}$ is finite and there exists an $H$-invariant partition $\Omega_{H,\infty}=\Omega_{1}\cup\cdots\cup\Omega_{k}$ into infinite subsets such that $H$ contains $\operatorname{Alt}_{f}(\Omega_{1})\times\cdots\times\operatorname{Alt}_{f}(\Omega_{k})$.
For $r\geq 2$, let $\Xi_{r}$ be the graph obtained by gluing $r$ infinite rays $\ell_{1},\cdots,\ell_{r}$ at their origin. The Houghton group $H_{r}$ is the group of all permutations $g$ of the vertex set of $\Xi_{r}$, such that there exists finite subset $E,F\subset\Xi_{r}$ such that for every $i$, $g|_{\ell_{i}\setminus E}$ is an isometry onto $\ell_{i}\setminus F$. The group $H_{r}$ is finitely generated for $r\geq 2$. Since the action of $H_{r}$ on $\Xi_{r}$ is by permutations of bounded displacements, it satisfies $\operatorname{vol}_{H_{r},\Xi_{r}}(n)\simeq n$.
In what follows, we fix $r\geq 2$. Let us denote $\Omega$ the vertex set of $\Xi_{r}$, and by $A:=\operatorname{Alt}_{f}(\Omega)$. Recall that $A$ is simple and contained in every non-trivial normal subgroup of $\operatorname{Sym}(\Omega)$. In particular $A$ is a normal subgroup of $H_{r}$ (and the quotient $H_{r}/A$ can easily be seen to be isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}^{r-1}$).
@@
Let $G:=H_{r}$ be the Houghton group, with $r\geq 2$. Let $X$ be a faithful $G$-set such that $\operatorname{vol}_{G,X}(n)\not\succeq n^{2}$. Let $X=\sqcup_{i\in I}X_{i}$ be the decomposition of $X$ into $G$-orbits. Then for every $i$ either the action of $G$ on $X_{i}$ is not faithful (equivalently, the group $A$ acts trivially), or it is conjugate to the diagonal action on $\Omega\times Y$, where $Y$ is a finite $G$-set whose cardinality is uniformly bounded (that is, the upper bound that does not depend on the choice of the orbit $X_{i}$).
@@
[…]
Let us first show that $|\Omega_{K,f}|\leq 1$. Suppose by contradiction that this is not the case. Let $K_{0}\leq K$ be the pointwise stabilizer of $\Omega_{K,f}$, which has finite index. Then $\operatorname{vol}_{G,G/K_{0}}(n)\simeq\operatorname{vol}_{G,X_{i}}(n)$ by Lemma 3.5. Choose two distinct points $x_{1},x_{2}\in\Omega_{K,f}$ and let $K_{1}\geq K_{0}$ be their pointwise stabiliser. Since $\operatorname{vol}_{G,X_{i}}(n)\succcurlyeq\operatorname{vol}_{G,G/K_{1}}(n)$, a contradiction will follow if we show that $\operatorname{vol}_{G,G/K_{1}}(n)\succcurlyeq n^{2}$. Since the action of $G$ on $\Omega$ is highly transitive (that is, transitive on $n$-tuples of distinct points), upon replacing $K_{1}$ by some $G$-conjugate we suppose that $x_{i}$ is the point at position 1 on $\ell_{i}$ for $i=1,2$ (with the convention that the common origin is at position 0).
@@
Note that $\lVert\gamma_{i,n}\rVert\leq Cn$, with $C=\max\{\lVert\sigma_{i}\rVert,\lVert t_{i}\rVert\;|\;i=1,2\}$. On the other hand each $\gamma_{i,n}$ is a permutation with finite support contained in $\ell_{i}$, and $\gamma_{i,n}(x_{i})$ is the point at position $n+1$ on $\ell_{i}$. Hence applying products of the form $\gamma_{1,m_{1}}\gamma_{2,m_{2}}$ to the pair $(x_{1},x_{2})$, with $0\leq m_{1},m_{2}\leq n$ shows that $\operatorname{vol}_{G,G/K_{1}}(n)\succcurlyeq n^{2}$. This is a contradiction and proves the claim that $|\Omega_{K,f}|\leq 1$.
@@
It follows that $K$ contains $\operatorname{Alt}(\Omega_{K,\infty})$. This also shows that $\Omega_{K,f}\neq\varnothing$, since otherwise $K$ would contain the normal subgroup $A$, contradicting that the action on $X_{I}$ is faithful. Hence $K\cap A$ contains a point stabiliser for the action of $A$ on $\Omega$, and hence is equal to it, as the latter is a maximal subgroup of $A$. Since $K$ was an arbitrary point stabiliser, we deduce that the action of $A$ on $X_{i}$ is conjugate on each of its orbits to the standard action of $A$ on $\Omega$. For each $A$-orbit $\mathcal{O}\subset X_{i}$ we denote by $j_{\mathcal{O}}\colon\mathcal{O}\to\Omega$ the unique $A$-equivariant bijection (its uniqueness follows from the fact that the only $A$-equivariant permutation of $\Omega$ is the identity).
Let $Y=X_{i}/A$, the space of $A$-orbits, on which $G$ acts since $A$ is normal. For a point $x\in A$, we denote by $\mathcal{O}(x)$ its $A$-orbit. Consider the map $\iota\colon X_{i}\to\Omega\times Y$ given by $\iota(x)=(j_{\mathcal{O}(x)}(x),\mathcal{O}(x))$. This map is equivariant (this is obvious for the second component; for the first, it follows from the observation that the map $g^{-1}\circ j_{\mathcal{O}(gx)}\circ g$ is an $A$-equivariant bijection from $\mathcal{O}(x)$ to $\Omega$, and thus it is equal to $j_{\mathcal{O}(x)}$). The map $\iota$ is also clearly injective. To check that it is surjective it is enough to observe that the diagonal $G$-action on $\Omega\times Y$ is transitive: this is true because the $G$-action on $Y$ is transitive, and the group $A$ acts trivially on $Y$ and transitively on each fiber $\Omega\times\{y\}$.
To conclude we need to argue that $Y$ must have uniformly bounded cardinality. Fix $n$ and suppose that the diameter of $Y$ is at least $n$. Then we can find a point $y\in Y$ and elements $s_{1},\ldots,s_{n}$ in the generating set of $G$ such that, if we set $g_{m}=s_{m}\cdots s_{1}$, then the points $y,g_{1}y,\ldots,g_{n}y$ are pairwise distinct (the point $y$ and sequence $(s_{i})$ can be found by looking at any path in the Schreier graph of $Y$ of length at least $n$). Next consider again the elements $\gamma_{i,n}$ above. Note that for every $n$, the element $\delta_{n}=\gamma_{1,n}\gamma_{2,n}$ belongs to $A$ and thus acts trivially on $Y$, and coincides with $\gamma_{1,n}$ in restriction to $\ell_{1}\subset\Omega$. Let $x\in\ell_{1}$ be at position 1, and consider the point $(x,y)\in\Omega\times Y$. Then the $n^{2}$ points $g_{m_{2}}\delta_{m_{1}}(x,y)=(g_{m_{2}}\gamma_{1,m_{1}}x,g_{m_{2}}y)$ for $m_{1},m_{2}\leq n$ are pairwise distinct, showing that there are at least $n^{2}$ in a ball of radius $O(n^{2})$. This can’t be true for $n$ arbitrarily large, showing that the cardinality of $Y$ must be uniformly bounded. ∎

### v2

@@ 3.1. A criterion
Let $G$ be a finitely generated group endowed with a finite symmetric generating set $S$. If $X$ is a $G$-set, we endow $X$ with its Schreier graph structure. The associated simplicial distance is denoted $d_{G,X}$ (we adopt the convention that $d_{G,X}(x,y)=+\infty$ for points in distinct $G$-orbits). We denote by $B_{G,X}(x,n)$ the corresponding balls. The support of an element $g\in G$ in $X$ is the set $\textrm{supp}_{X}(g)=\{x\in X\colon gx\neq x\}$. Given $R>0$, we shall say that two subsets $A,B\subset X$ are $R$-separated if $d_{G,X}(a,b)\geq R$ for every $a\in A$ and $b\in B$. Finally the $R$-coarse connected components of a subset $A\subset X$ are the equivalence classes of the equivalence relation on $A$ generated by the pairs $(a_{1},a_{2})$ such that $d_{G,X}(a_{1},a_{2})\leq R$.
@@
Let $G$ be a finitely generated group endowed with a finite symmetric generating set $S$. Let $\alpha,\beta\colon\mathbb{N}\to\mathbb{N}$ be functions such that $\alpha(n)\preccurlyeq\beta(n)$. We say that $G$ satisfies the sparse support condition at scale $(\alpha,\beta)$, if for every faithful $G$-set $X$ such that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\beta(n)$, there exist constants $C,D>0$ such that for every $R>0$, there exists a non-trivial element $g\in G$ satisfying the following:
@@
It is routine to check that this condition does not depend on the choice of a finite generating set $S$ for $G$ (using that a change of the generating set induces a bi-Lipschitz equivalence of the Schreier graphs, with constant depending on the generating sets only). Note that condition (C4) implies in particular that $G$ satisfies a Schreier growth gap $\alpha(n)$. The sparse support condition implies that this gap can be improved for any non-trivial free product with $G$, as follows.
@@
Set $L=G\ast H$. We fix a finite generating set of $L$ of the form $S\cup T$, where $S$ is a generating set of $G$ and $T$ is a generating set of $H$. Let $t\in T$ be a non-trivial generator of $H$. Let $X$ be a faithful $L$-set. We shall consider on $X$ the Schreier graph distance by $d_{L,X}$, as well as the distance $d_{G,X}$ induced by restricting the action to $G$. Suppose that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\beta(n)$ (else, the desired conclusion is true trivially). Let $C,D$ be as in Definition 3.1, fix $R>0$ (which we may assume is even), and let $g\in G$ be the corresponding element. Consider the commutator $h=[g,t]$. Note that $\textrm{supp}_{X}(h)$ is contained in the 1-neighbourhood of $\textrm{supp}_{X}(g)$ (with respect to the distance $d_{L,X}$).
Since $h$ has infinite order and the action is faithful, we can find $x_{0}\in X$ such that the points $x_{n}=h^{n}(x_{0})$ are pairwise distinct for $n=0,\ldots,R$. Note that $d_{L,X}(x_{n},x_{n+1})\leq 2D+2$, by condition (C1). For each point $x_{n}$ we choose $y_{n}\in\textrm{supp}_{X}(g)$ such that $d_{L,X}(x_{n},y_{n})\leq 1$. Then all points $y_{0},\ldots,y_{R}$ belong to the ball $B_{L,X}(x_{0},D_{1}R)$, with $D_{1}=2D+3$. Since all $D$-coarse connected components of $\textrm{supp}_{X}(g)$ with respect the distance $d_{G,X}$ have diameter bounded by $C$, their 1-neighbourhood in the distance $d_{L,X}$ has cardinality uniformly bounded by some constant $C_{1}>0$ (not depending on $R$).
@@
Remark 3.3.
We point out that in the proof of Proposition 3.2, the only property of the free product $L=G\ast H$ that has been used is the following: $L$ is an overgroup of $G$ containing an element $t\in L$ such that for every non-trivial $g\in G$, the commutator $[t,g]$ has infinite order.
@@ 3.2. Confined subgroups
Equivalently, a subgroup $H$ is confined if the closure of the set of conjugates of $H$ in the space $\operatorname{Sub}(G)$ of subgroups of $G$, does not contain the trivial subgroup. The following simple lemma explains the usefulness of the notion of confined subgroups for the study of growth of actions. See [LBMB22b, Lemma 1.8] for a proof. If $X$ is a $G$-set and $x\in X$, we denote by $G_{x}$ the stabilizer of $x$ in $G$.
@@
Here $\operatorname{vol}_{G}(n)$ represents the standard word growth given by the size of an $n$-ball in the Cayley graph of $G$ with respect some finite generating subset.
We will invoke the following, which is a particular case of Proposition 1.6 in [LBMB22b].
@@ 3.3. Houghton groups
If $\Omega$ is a set, $\operatorname{Sym}(\Omega)$ is the group of permutations of $\Omega$. We denote by $\operatorname{Alt}_{\mathrm{f}}(\Omega)$ the group of alternating finitely supported permutations of $\Omega$. For a subgroup $H\leq\operatorname{Sym}(\Omega)$, write $\Omega_{H,\mathrm{f}}$ for the union of finite orbits under $H$, and $\Omega_{H,\infty}$ for the union of infinite ones. We say that a partition is $H$-invariant if every element of $H$ sends a block of the partition to a (possibly different) block of the partition.
The following is a special case of [LBMB22a, Proposition 3.7].
@@
Let $\Omega$ be a countable set, and let $G\leq\operatorname{Sym}(\Omega)$ be any subgroup containing $\operatorname{Alt}_{\mathrm{f}}(\Omega)$. Then a subgroup $H\leq G$ is confined if and only if $\Omega_{H,\mathrm{f}}$ is finite and there exists an $H$-invariant partition $\Omega_{H,\infty}=\Omega_{1}\cup\cdots\cup\Omega_{k}$ such that $H$ contains $\operatorname{Alt}_{\mathrm{f}}(\Omega_{1})\times\cdots\times\operatorname{Alt}_{\mathrm{f}}(\Omega_{k})$.
For $r\geq 2$, let $\Xi_{r}$ be the graph obtained by gluing $r$ infinite rays $\ell_{1},\cdots,\ell_{r}$ at their origin. The Houghton group $H_{r}$ is the group of all permutations $g$ of the vertex set of $\Xi_{r}$, such that there exists finite subset $E,F\subset\Xi_{r}$ such that for every $i$, $g|_{\ell_{i}\setminus E}$ is an isometry onto $\ell_{i}\setminus F$. The group $H_{r}$ is finitely generated for $r\geq 2$. Since the action of $H_{r}$ on $\Xi_{r}$ is by permutations of bounded displacement, it satisfies $\operatorname{vol}_{H_{r},\Xi_{r}}(n)\simeq n$.
In what follows, we fix $r\geq 2$. Let us denote $\Omega$ the vertex set of $\Xi_{r}$, and by $A:=\operatorname{Alt}_{\mathrm{f}}(\Omega)$. Recall that $A$ is simple, and every non-trivial subgroup of $\operatorname{Sym}(\Omega)$ normalized by $A$ contains $A$. In particular $A$ is a normal subgroup of $H_{r}$.
@@
For $r\geq 2$, let $X$ be a faithful $H_{r}$-set such that $\operatorname{vol}_{G,X}(n)\not\succeq n^{2}$. Let $X=\sqcup_{i\in I}X_{i}$ be the decomposition of $X$ into $H_{r}$-orbits, and let $J$ be the subset of $i\in I$ such that the $H_{r}$-action on $X_{i}$ is faithful. Then:
  1. (1)
    for $i\notin J$, the $A$-action on $X_{i}$ is trivial;
  2. (2)
    there exist finite $H_{r}$-sets $(Y_{i})_{i\in J}$ of bounded cardinality (with respect to $i$) such that for every $i\in J$, the $H_{r}$-set $X_{i}$ is isomorphic to the product $H_{r}$-set $\Omega\times Y_{i}$.
In particular, if $Z=\bigcup_{i\notin J}X_{i}$ is the set of $A$-fixed points, there exists a finite index subgroup $K$ of $H_{r}$ such that every $K$-orbit in $X\setminus Z$ is isomorphic to $\Omega$ as a $K$-set.
@@
Set $G=H_{r}$. First note that item (1) follows from the fact that $A$ is contained in every non-trivial normal subgroup of $G$. We will show (2), which is the content of the statement. We fix a word metric $\lVert\cdot\rVert$ on $G$, associated to some symmetric generating set. Let $X$ be a $G$-set as in the statement and $K$ be the stabiliser of a point $x_{0}\in X_{i}$, with $I\in J$. The group $G$ has exponential growth, so by Lemma 3.5, $K$ must be confined. Thus, we can apply Proposition 3.7.
Let us first show that $|\Omega_{K,\mathrm{f}}|\leq 1$. Suppose by contradiction that this is not the case. Let $K_{0}\leq K$ be the pointwise stabilizer of the finite set $\Omega_{K,\mathrm{f}}$, which has finite index in $K$. Then $\operatorname{vol}_{G,G/K_{0}}(n)\simeq\operatorname{vol}_{G,X_{i}}(n)$ by Lemma 3.6. Choose two distinct points $x_{1},x_{2}\in\Omega_{K,\mathrm{f}}$ and let $K_{1}\geq K_{0}$ be their pointwise stabiliser. Since $\operatorname{vol}_{G,X_{i}}(n)\succcurlyeq\operatorname{vol}_{G,G/K_{1}}(n)$, a contradiction will follow if we show that $\operatorname{vol}_{G,G/K_{1}}(n)\succcurlyeq n^{2}$. Since the action of $G$ on $\Omega$ is highly transitive (that is, transitive on $n$-tuples of distinct points), upon replacing $K_{1}$ by some $G$-conjugate we suppose that $x_{i}$ is the point at position 1 on $\ell_{i}$ for $i=1,2$ (with the convention that the common origin is at position 0).
@@
Note that $\lVert\gamma_{i,n}\rVert\leq Cn$, with $C=\max\{\lVert\sigma_{i}\rVert,\lVert t_{i}\rVert\;|\;i=1,2\}$. On the other hand each $\gamma_{i,n}$ is a permutation with finite support contained in $\ell_{i}$, and $\gamma_{i,n}(x_{i})$ is the point at position $n+1$ on $\ell_{i}$. Hence applying products of the form $\gamma_{1,m_{1}}\gamma_{2,m_{2}}$ to the pair $(x_{1},x_{2})$, with $0\leq m_{1},m_{2}\leq n$ shows that $\operatorname{vol}_{G,G/K_{1}}(n)\succcurlyeq n^{2}$. This is a contradiction and proves the claim that $|\Omega_{K,\mathrm{f}}|\leq 1$.
@@
It follows that $K$ contains $\operatorname{Alt}(\Omega_{K,\infty})$. This also shows that $\Omega_{K,\mathrm{f}}\neq\varnothing$, since otherwise $K$ would contain the normal subgroup $A$, contradicting that the action on $X_{I}$ is faithful. Hence $K\cap A$ contains a point stabiliser for the action of $A$ on $\Omega$, and hence is equal to it, as the latter is a maximal subgroup of $A$. Since $K$ was an arbitrary point stabiliser, we deduce that the action of $A$ on $X_{i}$ is conjugate on each of its orbits to the standard action of $A$ on $\Omega$. For each $A$-orbit $\mathcal{O}\subset X_{i}$ we denote by $j_{\mathcal{O}}\colon\mathcal{O}\to\Omega$ the unique $A$-equivariant bijection (its uniqueness follows from the fact that the only $A$-equivariant permutation of $\Omega$ is the identity).
Let $Y_{i}=X_{i}/A$, the space of $A$-orbits, on which $G$ acts since $A$ is normal. For a point $x\in A$, we denote by $\mathcal{O}(x)$ its $A$-orbit. Consider the map $\iota\colon X_{i}\to\Omega\times Y_{i}$ given by $\iota(x)=(j_{\mathcal{O}(x)}(x),\mathcal{O}(x))$. This map is equivariant (this is obvious for the second component; for the first, it follows from the observation that the map $g^{-1}\circ j_{\mathcal{O}(gx)}\circ g$ is an $A$-equivariant bijection from $\mathcal{O}(x)$ to $\Omega$, and thus it is equal to $j_{\mathcal{O}(x)}$). The map $\iota$ is also clearly injective. To check that it is surjective it is enough to observe that the diagonal $G$-action on $\Omega\times Y_{i}$ is transitive: this is true because the $G$-action on $Y_{i}$ is transitive, and the group $A$ acts trivially on $Y_{i}$ and transitively on each fiber $\Omega\times\{y\}$.
To conclude we need to argue that $Y_{i}$ must have uniformly bounded cardinality when $i\in J$ varies. Fix $n$ and suppose that the diameter of $Y_{i}$ is at least $n$. Then we can find a point $y\in Y_{i}$ and elements $s_{1},\ldots,s_{n}$ in the generating set of $G$ such that, if we set $g_{m}=s_{m}\cdots s_{1}$, then the points $y,g_{1}y,\ldots,g_{n}y$ are pairwise distinct (the point $y$ and sequence $(s_{i})$ can be found by looking at any path in the Schreier graph of $Y_{i}$ of length at least $n$). Next consider again the elements $\gamma_{i,n}$ above. Note that for every $n$, the element $\delta_{n}=\gamma_{1,n}\gamma_{2,n}$ belongs to $A$ and thus acts trivially on $Y_{i}$, and coincides with $\gamma_{1,n}$ in restriction to $\ell_{1}\subset\Omega$. Let $x\in\ell_{1}$ be at position 1, and consider the point $(x,y)\in\Omega\times Y_{i}$. Then the $n^{2}$ points $g_{m_{2}}\delta_{m_{1}}(x,y)=(g_{m_{2}}\gamma_{1,m_{1}}x,g_{m_{2}}y)$ for $m_{1},m_{2}\leq n$ are pairwise distinct, showing that there are at least $n^{2}$ in a ball of radius $O(n^{2})$. This cannot be true for $n$ arbitrarily large, showing that the cardinality of $Y_{i}$ must be uniformly bounded.
Finally, since the group $G$ is finitely generated, we may choose a finite index subgroup $K$ which acts trivially on $Y_{i}$ for every $i$, showing the last sentence in the statement. ∎
@@
Remark 3.10.
We point out that the proof of Theorem 3.9, the assumption that the group $L:=H_{r}\ast K$ is a free product is only used to invoke Proposition 3.2. As a consequence, this assumption can be relaxed as in Remark 3.3.
Remark 3.11.
To prove Theorem 3.9 we could have restricted to the case $r=2$, since every $H_{r}$ contains $H_{2}$ as a subgroup. Note however that the previous proof provides additional information on actions of small growth of $H_{r}$ for general $r$ (Proposition 3.8).